OpenStudy (anonymous):

The sequence is geometric. What is the next term? 12, 24, 48, 96,... 186 192 288

OpenStudy (anonymous):

192

OpenStudy (kinggeorge):

Notice that \(24=2*12\), \(48=2*24\), and that \(96=2*48\). If we follow this pattern, what do you think will come next?

OpenStudy (anonymous):

192?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

Precisely. \[192=2*96\]So 192 would be the next term.